# Quantum-Enhanced Power Flow and Optimal Power Flow based on Combinatorial Reformulation

### v2

@@ Abstract
This study introduces the Adiabatic Quantum Power Flow (AQPF) and Adiabatic Quantum Optimal Power Flow (AQOPF) algorithms to solve power flow (PF) and optimal power flow (OPF) problems, respectively. These algorithms utilize a novel combinatorial optimization reformulation of classical PF and OPF problems, and hence, enable their implementation on Ising machines, e.g., quantum and quantum-inspired hardware. The experiments are conducted on standard test cases ranging from 4-bus to 1354-bus systems, using D-Wave’s Advantage™  system (QA), its hybrid quantum-classical solver (HA), as well as the third-generation Digital Annealer (DAv3) and Quantum-Inspired Integrated Optimization software (QIIO) developed by Fujitsu. The annealers are systematically evaluated based on: (i) full and partitioned formulations, (ii) ability to handle ill-conditioned cases, and (iii) scalability. The results are benchmarked against the Newton-Raphson numerical method (NR) and suggest that AQPF and AQOPF can serve as effective solvers or complementary tools to classical methods to address unsolved challenges in large-scale modern power systems.
Index Terms:
Adiabatic quantum computing, hybrid quantum-classical solvers, quantum annealing, noisy intermediate-scale quantum (NISQ) era, state estimation.
I Introduction
Power flow (PF) and optimal power flow (OPF) are fundamental tasks to power system operation, planning, and control [1]. PF analysis determines the steady-state voltages, power injections, and power flows in electrical grids under given operating conditions. OPF extends PF analysis by minimizing generation costs, reducing power losses, and/or improving voltage stability while satisfying operational constraints [2]. With the increasing integration of renewable energy sources, demand-side management, and network congestion challenges, efficient OPF solutions are critical for ensuring grid reliability (e.g., [3]).
@@ I Introduction
The Newton-Raphson (NR) method is widely used to solve PF and OPF problems due to its fast convergence rate and robustness for well-conditioned cases [4]. NR is an iterative numerical method that linearizes the PF equations using the Jacobian matrix and updates the solution until convergence is achieved. However, despite its efficiency in small- to medium-sized systems, NR faces significant computational challenges when applied to large-scale electrical grids with thousands of buses [5]. These challenges arise from repeated Jacobian matrix computations, matrix inversions, and high memory requirements, which lead to increased computational time and convergence issues. In addition, NR struggles with ill-conditioned power systems scenarios, where numerical instability occurs due to poorly scaled Jacobian matrices. Such cases often arise in networks with high R/X ratios, weakly connected areas, or stressed operating conditions, for instance, during contingencies, voltage collapse scenarios, or low-inertia grid conditions [6]. These limitations make NR impractical for real-time power system operation, motivating the search for alternative methods [7].
@@
Several approaches have been proposed in the literature to address the limitations of NR for large-scale and ill-conditioned power systems. For example, a fast NR-based PF algorithm is developed in [8] that uses sparse matrix techniques and parallel processing to improve computational efficiency, particularly for large-scale power systems. Another example is [9] where the authors propose a hybrid approach that integrates NR with stochastic gradient descent to enhance convergence properties, especially in ill-conditioned cases. In [10], a fully parallel NR implementation is also introduced that utilizes GPU-CPU vectorization and sparse techniques to accelerate PF analysis. Finally, an alternative PF controller is explored in [11] to improve PF capability, indirectly addressing some of the limitations of conventional NR-based methods. Note that while parallel NR implementations improve scalability, they still rely on matrix operations that become increasingly expensive as network sizes grow. Machine learning-based approaches, on the other hand, have been proposed to approximate PF and OPF solutions, but their accuracy depends on training data, limiting their generalizability (e.g., [12, 13, 14]). These studies demonstrate that enhancing NR with advanced numerical techniques, hybrid approaches, and parallelization can significantly improve its applicability to large-scale power systems, yet fundamental challenges remain in terms of computational complexity and convergence reliability.
@@
Given these computational challenges, there is growing interest in quantum computing to address PF and OPF problems. Ising machines, in general, and quantum/digital annealing, in particular, offer the ability to efficiently explore large solution spaces, which makes them a promising candidate for tackling combinatorial optimization problems in power systems [15]. In contrast to classical iterative solvers, quantum approaches aim to directly minimize energy functions associated with PF and OPF formulations, potentially achieving superior scalability and convergence [16]. This work extends the Adiabatic Quantum Power Flow (AQPF) algorithm, proposed in [16], to solve OPF problems, introducing the Adiabatic Quantum Optimal Power Flow (AQOPF) algorithm. AQPF and AQOPF utilize a novel combinatorial optimization formulation that reformulates and restructures classical PF and OPF problems, and hence enables their implementation on quantum and quantum-inspired hardware. The experiments are conducted on standard test cases ranging from 4-bus to 1354-bus systems, using D-Wave’s Advantage™  system (QA), its hybrid quantum-classical solver (HA), as well as the third-generation Digital Annealer (DAv3) and Quantum-Inspired Integrated Optimization software (QIIO) developed by Fujitsu. The key contributions of this paper are:
@@
  • The AQPF algorithm, proposed in [16], is extended to solve OPF problems, introducing the AQOPF algorithm. AQPF is also evaluated using the latest Fujitsu’s QIIO software to assess its performance on a state-of-the-art quantum-inspired solver.
@@
  • AQPF and AQOPF are shown to effectively handle ill-conditioned cases where conventional solvers fail to converge. The partitioned formulation also exhibits similar robustness in handling such cases.
  • AQPF and AQOPF are shown to be capable of addressing large-scale systems, provided the solver can handle a high number of binary variables.
II Mathematical formulation
@@
where $S_{i}$ denotes the net power flowing through bus $i$, $S_{i}^{G}$ and $S_{i}^{D}$ represent the total power generated and total power demand at bus $i$, respectively. Further expanding 2 into its real and imaginary components gives:
$P_{i}+jQ_{i}=(P_{i}^{G}-P_{i}^{D})+j(Q_{i}^{G}-Q_{i}^{D}),\quad i\in\{1,\dots,N\},$ (3)
@@
Further expanding 4 into real and imaginary components yields the AC PF equations in rectangular coordinates:
@@
Here, the complex voltage at bus $i\in\{1,\dots,N\}$ is expressed as ${V}_{i}=\mu_{i}+j\omega_{i}$, with $\mu_{i}$ and $\omega_{i}$ respectively representing its real and imaginary parts.
II-A Power Flow Formulation
The objective of power flow (PF) analysis is to compute the complex bus voltages and power injections that satisfy the power balance equations, as given in:
@@ II-A Power Flow Formulation
These equations are iteratively solved, typically using numerical methods, such as the Newton-Raphson (NR) or Gauss-Seidel algorithms, to determine the complex bus voltages, i.e., $\mu_{i}$ and $\omega_{i}$ in rectangular form such that equations 6 hold for all buses $i\in\{1,\dots,N\}$.
II-B Optimal Power Flow Formulation
The OPF problem aims to minimize total generation costs while ensuring that all system constraints are satisfied. That is, the goal is to compute the optimal generator outputs $\{P_{i}^{G},Q_{i}^{G}\}$ and bus voltages $\{V_{i},\delta_{i}\}$ that minimize generation costs while satisfying the power balance, generation, and operational limits. The problem can be formulated as:
@@ II-B Optimal Power Flow Formulation
$\underline{\delta_{i}}$ $\leq\delta_{i}\leq\overline{\delta_{i}}.$ (8f)
@@ III Combinatorial formulation
III-A Combinatorial Power Flow Formulation
@@
where the goal is to minimize this mismatch to satisfy the power balance equations (6). To further convert the problem into a form that annealers can handle, we expand 5 into:
@@
Here, the variables $\mu_{i}$, $\mu_{k}$, $\omega_{i}$, and $\omega_{k}$ continuous real numbers that must be discretized to reformulate the problem as a binary combinatorial optimization task. This leads naturally to a Quadratic Unconstrained Binary Optimization (QUBO) model, which is defined by its symmetric, real-valued matrix $Q\in\mathbb{R}^{n\times n}$ and the binary minimization problem:
$\min_{\mathbf{x}\in\{0,1\}^{n}}\sum_{i=1}^{n}\sum_{j=1}^{n}Q_{ij}x_{i}x_{j},$ (11)
@@
By inserting 12 into 10a, we derive an expression for the net active power injection $P_{i}$. Note that this expression follows the derivation in our previous work [16], and is reproduced here for completeness:
$\begin{split}P_{i}=\sum_{j=0}^{N}\Big{[}\mu_{i}^{0}G_{ij}\mu_{j}^{0}+\omega_{i}^{0}G_{ij}\omega_{j}^{0}+\omega_{i}^{0}B_{ij}\mu_{j}^{0}-\mu_{i}^{0}B_{ij}\omega_{j}^{0}\Big{]}\\ +\sum_{j=0}^{N}\sum_{k=0}^{1}\Big{[}(-1)^{k}\mu_{i}^{0}G_{ij}x_{j,k}^{\mu}\Delta\mu_{j}+(-1)^{k}x_{i,k}^{\mu}\Delta\mu_{i}G_{ij}\mu_{j}^{0}\\ +(-1)^{k}\omega_{i}^{0}G_{ij}x_{j,k}^{\omega}\Delta\omega_{j}+(-1)^{k}x_{i,k}^{\omega}\Delta\omega_{i}G_{ij}\omega_{j}^{0}\\[8.61108pt] +(-1)^{k}\omega_{i}^{0}B_{ij}x_{j,k}^{\mu}\Delta\mu_{j}+(-1)^{k}x_{i,k}^{\omega}\Delta\omega_{i}B_{ij}\mu_{j}^{0}\\[8.61108pt] -(-1)^{k}\mu_{i}^{0}B_{ij}x_{j,k}^{\omega}\Delta\omega_{j}-(-1)^{k}x_{i,k}^{\mu}\Delta\mu_{i}B_{ij}\omega_{j}^{0}\Big{]}\\[8.61108pt] +\sum_{j=0}^{N}\sum_{k=0}^{1}\sum_{l=0}^{1}\Big{[}(-1)^{k+l}x_{i,k}^{\mu}\Delta\mu_{i}G_{ij}x_{j,l}^{\mu}\Delta\mu_{j}\\[8.61108pt] +(-1)^{k+l}x_{i,k}^{\omega}\Delta\omega_{i}G_{ij}x_{j,l}^{\omega}\Delta\omega_{j}\\[8.61108pt] +(-1)^{k+l}x_{i,k}^{\omega}\Delta\omega_{i}B_{ij}x_{j,l}^{\mu}\Delta\mu_{j}\\[8.61108pt] -(-1)^{k+l}x_{i,k}^{\mu}\Delta\mu_{i}B_{ij}x_{j,l}^{\omega}\Delta\omega_{j}\Big{]},\end{split}$ (13)
An analogous expression for the net reactive power $Q_{i}$ is obtained by inserting (12) into (10b). This expression is also based on our earlier work [16]:
@@
The ansatz presented in 13 and 14 yields a QUBO structure, as expressed in (11), enabling the minimization of the squared mismatch terms in 9 using annealing approaches:
[…]
@@
III-B Combinatorial Optimal Power Flow Formulation
In contrast to the previously described PF analysis, power balance equations (6) become equality constraints in the optimal PF problem, i.e., 8a and 8b, and, consequently, (15) becomes a penalty term in the QUBO formulation of the OPF problem, $\lambda_{\text{PF}}H_{\text{obj}}(\cdot)$, where $\lambda_{\text{PF}}$ is the corresponding penalty parameter. Another penalty term is needed to implement the inequality constraint terms, $H_{\text{const}}(\cdot)$, as well as the objective function of the OPF problem as an equality constraint term, $H_{\text{cost}}(\cdot)$.
Let us start by describing the Lagrange multiplier approach, which is used to impose equality constraints in a weak sense in a QUBO formulation. To this end, consider the following minimization problem with a single equality constraint:

### v3

@@ Abstract
This study further explores reformulating power flow (PF) analysis as a discrete combinatorial optimization problem, proposed in our earlier study using the Adiabatic Quantum Power Flow (AQPF) algorithm, which can be executed on Ising machines, including quantum and quantum-inspired hardware. This approach provides a new representation of the underlying equations, analogous to how neural networks approximate complex functions using simple operations. While the resulting combinatorial optimization problem is NP-hard, it is compatible with emerging quantum hardware designed to address such complexity. We introduce the Adiabatic Quantum Optimal Power Flow (AQOPF) algorithm, which transforms the classical optimal power flow (OPF) equations into quadratic unconstrained binary optimization (QUBO) models. Furthermore, the AQPF and AQOPF algorithms are evaluated on standard test cases ranging from 4- to 1354-bus systems using D-Wave’s Advantage™ system (QA), its hybrid quantum-classical solver (HA), and Fujitsu’s third-generation Digital Annealer (DAv3) and Quantum-Inspired Integrated Optimization (QIIO) platform. Both full and partitioned formulations are investigated, with particular attention to scalability and robustness in ill-conditioned scenarios. The results demonstrate that the algorithms can reproduce feasible PF and OPF solutions and exhibit promising computational scalability when supported by scalable hardware.
keywords:
[…]
1 Introduction
@@
Power flow (PF) and optimal power flow (OPF) are fundamental tasks in power system operation, planning, and control. PF analysis determines the steady-state voltages, power injections, and power flows in electrical grids under given operating conditions. OPF extends PF analysis by minimizing generation costs, reducing power losses, and/or improving voltage stability while satisfying operational constraints Xiang2026 ; Jing2023 . With the increasing integration of renewable energy sources, demand-side management, and network congestion, efficient OPF solutions are critical to ensuring grid reliability Li2025 . The Newton-Raphson (NR) method is widely used to solve PF and OPF problems due to its fast convergence rate and robustness for well-conditioned cases. It is an iterative numerical method that linearizes the PF equations using the Jacobian matrix and updates the solution until convergence is achieved. However, despite its efficiency in small- to medium-sized systems, NR faces significant computational challenges when applied to large-scale electrical grids with thousands of buses Liu2022 . These challenges arise from repeated Jacobian matrix computations, matrix inversions, and high memory requirements, resulting in increased computational time and convergence issues. In addition, NR struggles with ill-conditioned power system scenarios, where numerical instability arises from poorly scaled Jacobian matrices. Such cases often arise in networks with high R/X ratios, weakly connected areas, or stressed operating conditions, for instance, during contingencies, voltage collapse scenarios, or low-inertia grid conditions Nur2021 . These limitations make NR impractical for real-time operation, motivating the search for alternative methods Liu2020_ .
@@
Several approaches have been proposed in the literature to address the limitations of NR in large-scale, ill-conditioned power systems. For example, a fast NR-based PF algorithm is developed in Ahmadi2022 that uses sparse matrix techniques and parallel processing to improve computational efficiency, particularly for large-scale power systems. Another example is Costilla-Enriquez2021 where the authors propose a hybrid approach that integrates NR with stochastic gradient descent to enhance convergence properties, especially in ill-conditioned cases. In Su2020 , a fully parallel NR implementation is also introduced that utilizes GPU-CPU vectorization and sparse techniques to accelerate PF analysis. Finally, an alternative PF controller is explored in  Kumari2020 to improve PF capability, indirectly addressing some of the limitations of conventional NR-based methods. Note that while parallel NR implementations improve scalability, they still rely on matrix operations that become increasingly expensive as network sizes grow. Machine learning-based approaches, on the other hand, have been proposed to approximate PF and OPF solutions, but their accuracy depends on training data, limiting their generalization ability Kaseb2026pinn ; Liu2024 ; Wu2024 . These studies demonstrate that enhancing NR with advanced numerical techniques, hybrid approaches, and parallelization can significantly improve its applicability to large-scale power systems. Yet, fundamental challenges remain regarding computational complexity and convergence reliability. Thus, there is growing interest in the emerging field of quantum computing to address power system problems. Ising machines, in general, and quantum/digital annealing, in particular, offer the ability to efficiently explore large solution spaces, which makes them a promising candidate for tackling combinatorial optimization problems in power systems Yin2023 .
It is, however, important to emphasize that the aim of this study is not to replace the highly efficient classical algorithms developed for PF and OPF over several decades, e.g., Su2020 ; Ahmadi2022 . Classical numerical solvers remain the most practical and reliable tools for PF and OPF in current power system applications. Instead, this study further explores the alternative representation of the PF and OPF equations through a discrete combinatorial optimization approach proposed in our earlier study kaseb2024power . Similar to how neural networks approximate complex mathematical relationships using simple algebraic operations, the proposed reformulation discretizes the governing equations. This alternative representation provides a different computational perspective on PF and OPF, and can potentially address challenges encountered by numerical solvers in ill-conditioned scenarios. However, this representation introduces NP-hard optimization problems, which are generally challenging for classical algorithms but are compatible with emerging quantum hardware Dent2026 . Although current quantum hardware remains limited in scale and precision, the development of algorithms that can operate on such devices is an important step toward understanding their potential future role in PF and OPF. The goal of this study is therefore to investigate whether PF and OPF can be reformulated in a way that may eventually enable quantum hardware to assist in addressing challenging scenarios that remain difficult for classical solvers, e.g., ill-conditioned networks.
@@
That said, the present study extends the Adiabatic Quantum Power Flow (AQPF) algorithm, proposed in kaseb2024power , to solve OPF problems, introducing the Adiabatic Quantum Optimal Power Flow (AQOPF) algorithm. The experiments are conducted on standard test cases ranging from 4- to 1354-bus systems, using D-Wave’s Advantage™  system (QA), its hybrid quantum-classical solver (HA), as well as the third-generation Digital Annealer (DAv3) and Quantum-Inspired Integrated Optimization software (QIIO) developed by Fujitsu. The contributions of this paper are:
@@
  • The AQPF algorithm, proposed in kaseb2024power , is extended to solve OPF problems, introducing the AQOPF algorithm. AQPF is also evaluated using Fujitsu’s latest QIIO software to assess its performance on a state-of-the-art quantum-inspired solver.
@@
  • AQPF and AQOPF are shown to handle ill-conditioned cases where conventional solvers fail to converge effectively. The partitioned formulation also exhibits similar robustness in handling such cases.
  • AQPF and AQOPF are shown to be computationally capable of addressing large-scale systems, provided the solver can handle a high number of binary variables.
2 Mathematical formulation
@@
where $S_{i}$ denotes the net power flowing through bus $i$, $S_{i}^{G}$ and $S_{i}^{D}$ represent the total power generated and total power demand at bus $i$, respectively. Further expanding ˜2 into its real and imaginary components gives:
$P_{i}+jQ_{i}=(P_{i}^{G}-P_{i}^{D})+j(Q_{i}^{G}-Q_{i}^{D}),\quad i\in\{1,\dots,N\}.$ (3)
@@
Further expanding ˜4 into real and imaginary components yields the AC PF equations in rectangular coordinates:
@@
where the complex voltage at bus $i\in\{1,\dots,N\}$ is expressed as ${V}_{i}=\mu_{i}+j\omega_{i}$, with $\mu_{i}$ and $\omega_{i}$ respectively representing its real and imaginary parts.
Each bus $i$ in a power system is defined by four primary variables: net active power injection ($P_{i}$), net reactive power injection ($Q_{i}$), voltage magnitude ($V_{i}$), and voltage angle ($\delta_{i}$). Based on the variables specified or controlled at each bus, buses in a power system are divided into three main types: PV bus, PQ bus, and slack bus. Table˜1 summarizes the characteristics of these bus types. The slack bus serves as a reference point for computing the PF and OPF.
Table 1: Types of buses in a power system.
slack bus PV bus PQ bus
Known variables $V_{i}$, $\delta_{i}$ $P_{i}$, $V_{i}$ $P_{i}$, $Q_{i}$
Unknown variables $P_{i}$, $Q_{i}$ $Q_{i}$, $\delta_{i}$ $V_{i}$, $\delta_{i}$
Number of buses in a network 1 $_{G} NN$$N_{G}-1$
2.1 Power Flow Formulation
The objective of PF analysis is to compute the complex bus voltages and power injections that satisfy the power balance equations, as given in:
@@ 2.1 Power Flow Formulation
These equations are iteratively solved, typically using numerical methods, such as the NR or Gauss-Seidel methods, to determine the complex bus voltages, i.e., $\mu_{i}$ and $\omega_{i}$ in rectangular form such that equations ˜6 hold for all buses $i\in\{1,\dots,N\}$.
2.2 Optimal Power Flow Formulation
The OPF problem aims to minimize total generation costs while ensuring that all system constraints are satisfied. The goal is to compute the optimal generator outputs $\{P_{i}^{G},Q_{i}^{G}\}$ and bus voltages $\{V_{i},\delta_{i}\}$ that minimize generation costs while satisfying the power balance, generation, and operational limits. The problem is formulated as:
@@ 2.2 Optimal Power Flow Formulation
$\underline{\delta_{i}}$ $\leq\delta_{i}\leq\overline{\delta_{i}},$ (8f)
@@ 3 Combinatorial formulation
3.1 Combinatorial Power Flow Formulation
@@
where the goal is to minimize this mismatch to satisfy the power balance equations (6). To further convert the problem into a form that annealers can handle, we expand ˜5 into:
@@
Here, the variables $\mu_{i}$, $\mu_{k}$, $\omega_{i}$, and $\omega_{k}$ are continuous real numbers that must be discretized to reformulate the problem as a binary combinatorial optimization task. This leads naturally to a Quadratic Unconstrained Binary Optimization (QUBO) model, which is defined by its symmetric, real-valued matrix $Q\in\mathbb{R}^{n\times n}$ and the binary minimization problem:
$\min_{\mathbf{x}\in\{0,1\}^{n}}\sum_{i=1}^{n}\sum_{j=1}^{n}Q_{ij}x_{i}x_{j}.$ (11)
@@
By inserting ˜12 into ˜10a, we derive an expression for the net active power injection $P_{i}$. Note that this expression follows the derivation in our previous work kaseb2024power , and is reproduced here for completeness:
$\begin{split}P_{i}=\sum_{j=0}^{N}\Big[\mu_{i}^{0}G_{ij}\mu_{j}^{0}+\omega_{i}^{0}G_{ij}\omega_{j}^{0}+\omega_{i}^{0}B_{ij}\mu_{j}^{0}-\mu_{i}^{0}B_{ij}\omega_{j}^{0}\Big]\\ +\sum_{j=0}^{N}\sum_{k=0}^{1}\Big[(-1)^{k}\mu_{i}^{0}G_{ij}x_{j,k}^{\mu}\Delta\mu_{j}+(-1)^{k}x_{i,k}^{\mu}\Delta\mu_{i}G_{ij}\mu_{j}^{0}\\ +(-1)^{k}\omega_{i}^{0}G_{ij}x_{j,k}^{\omega}\Delta\omega_{j}+(-1)^{k}x_{i,k}^{\omega}\Delta\omega_{i}G_{ij}\omega_{j}^{0}\\[8.61108pt] +(-1)^{k}\omega_{i}^{0}B_{ij}x_{j,k}^{\mu}\Delta\mu_{j}+(-1)^{k}x_{i,k}^{\omega}\Delta\omega_{i}B_{ij}\mu_{j}^{0}\\[8.61108pt] -(-1)^{k}\mu_{i}^{0}B_{ij}x_{j,k}^{\omega}\Delta\omega_{j}-(-1)^{k}x_{i,k}^{\mu}\Delta\mu_{i}B_{ij}\omega_{j}^{0}\Big]\\[8.61108pt] +\sum_{j=0}^{N}\sum_{k=0}^{1}\sum_{l=0}^{1}\Big[(-1)^{k+l}x_{i,k}^{\mu}\Delta\mu_{i}G_{ij}x_{j,l}^{\mu}\Delta\mu_{j}\\[8.61108pt] +(-1)^{k+l}x_{i,k}^{\omega}\Delta\omega_{i}G_{ij}x_{j,l}^{\omega}\Delta\omega_{j}\\[8.61108pt] +(-1)^{k+l}x_{i,k}^{\omega}\Delta\omega_{i}B_{ij}x_{j,l}^{\mu}\Delta\mu_{j}\\[8.61108pt] -(-1)^{k+l}x_{i,k}^{\mu}\Delta\mu_{i}B_{ij}x_{j,l}^{\omega}\Delta\omega_{j}\Big].\end{split}$ (13)
An analogous expression for the net reactive power $Q_{i}$ is obtained by inserting (12) into (10b). This expression is also based on our earlier work kaseb2024power :
@@
The ansatz presented in ˜13 and 14 yields a QUBO structure, as expressed in (11), enabling the minimization of the squared mismatch terms in ˜9 using annealing approaches:
$H_{\text{obj}}(\mathbf{x})=\sum\limits_{i=1}^{N}(P_{i}-P_{i}^{G}+P_{i}^{D})^{2}+(Q_{i}-Q_{i}^{G}+Q_{i}^{D})^{2},$ (15)
@@
3.2 Combinatorial Optimal Power Flow Formulation
In contrast to the previously described PF analysis, power balance equations (6) become equality constraints in the optimal PF problem, i.e., ˜8a and 8b, and, consequently, (15) becomes a penalty term in the QUBO formulation of the OPF problem, $\lambda_{\text{PF}}H_{\text{obj}}(\cdot)$, where $\lambda_{\text{PF}}$ is the corresponding penalty parameter. Another penalty term is needed to implement the inequality constraint terms, $H_{\text{const}}(\cdot)$, as well as the objective function of the OPF problem as an equality constraint term, $H_{\text{cost}}(\cdot)$. Let us start by describing the Lagrange multiplier approach, which imposes equality constraints in a weak sense within a QUBO formulation. To this end, consider the following minimization problem with a single equality constraint: